OpenStudy (agent_a):

Use the given parameters to answer the following questions. x = cos(3θ) y = 5sin(θ) (a) Find the points on the curve where the tangent is horizontal. (b) Find the points on the curve where the tangent is vertical. I have the answers, but for part b, 2 out of my 6 answers are wrong. Can you help me out? Thanks in advance!

OpenStudy (agent_a):

The equations: \[x = \cos(3\theta)\] \[y=5\sin(\theta)\]

OpenStudy (dumbcow):

well slope of tangent line is dy/dx \[\frac{dy}{dx} = \frac{dy}{d \theta}*\frac{d \theta}{dx}\] \[= 5\cos \theta * \frac{1}{-3 \sin (3 \theta)} = -\frac{5}{3} \frac{\cos \theta}{\sin (3 \theta)}\] tangent is horizontal when slope = 0 \[\rightarrow -\frac{5 \cos \theta}{3 \sin(3 \theta)} = 0\] this occurs when cos = 0 \[\theta = \frac{\pi}{2} + k \pi\] plugging this in you get points (0,5) and (0,-5) tangent is vertical when slope is infinite or undefined this occurs when denominator approaches zero \[\sin(3 \theta) = 0\] \[3 \theta = k \pi\] \[\theta =\frac{k \pi}{3}\] plugging this in you get points (1,0) , (-1, 5sqrt3/2) , (1, 5sqrt3/2) , (-1,0), (1, -5sqrt3/2) and (-1, -5sqrt3/2)

OpenStudy (agent_a):

Thank You!